OpenStudy (erinweeks):

Help Me (; . Salisbury High School decided to take their students on a field trip to a theme park. A total of 150 people went on the trip. Adults pay $45.00 for a ticket and students pay $28.50 for a ticket. How many students and how many adults went to the park if they paid a total of $4770? (2 points)

OpenStudy (merchandize):

A+S=150 OR A=150-S 45A+28.50S=4,770 45(150-S)+28.50S=4,770 6,750-45S+28.50S=4,770 -16.50S=4,770-6,750 -16.50S=-1,980 S=-1,980/-16.50 S=120 STUDENT TICKETS WERE SOLD. 150-120=30 ADULT TICKETS WERE SOLD. PROOF: 45*30+28.50*120=4,770 1,350+3,420=4,770 4,770=4,770

OpenStudy (erinweeks):

Good looks on the help you gave me I really appreciate ;)

OpenStudy (merchandize):

thanks